During a science experiment, Alma created a 22% alcohol solution by mixing 8 fluid ounces of a 28% alcohol solution with a certain amount of a 6% alcohol solution. How many fluid ounces of the 22% alcohol solution did she create?
Answer by rothauserc(4718)

let x be the amount of the 22% solution
let y be the amount of the 6% solution
:
We start with what we know
:
8 * 0.28 = 2.24
y * 0.06 = 0.06y
x * 0.22 = 0.22x
:
x = y + 8
:
y = x - 8
:
This equation represents mixing the 6% solution with the 28% solution
:
0.06(x-8) +2.24 = 0.22x
:
0.06x -0.48 +2.24 = 0.22x
:
0.16x = 1.76
:
x = 11
:
******************************************
Alma created 11 ounces of the 22% solution
******************************************
: